OpenStudy (jacobbirddog):

Please Help!!!!! Fan+medal! If x varies inversely with y and x = 6 when y = 10, find y when x = 8.

OpenStudy (anonymous):

how can i help you

OpenStudy (jacobbirddog):

well i need the answer in fraction formand i'm wondering how would i get it?

OpenStudy (anonymous):

luwigi

OpenStudy (anonymous):

sorry

OpenStudy (jacobbirddog):

ok?

OpenStudy (anonymous):

what is the question

OpenStudy (jacobbirddog):

If x varies inversely with y and x = 6 when y = 10, find y when x = 8.

OpenStudy (anonymous):

it does not make sense to me sorry

ganeshie8 (ganeshie8):

Inverse relation : \(\large y = \frac{k}{x} \)

ganeshie8 (ganeshie8):

plugin x = 6, y = 10 and find the value of \(\large k\) first

OpenStudy (jacobbirddog):

ok

OpenStudy (jacobbirddog):

would it be 6/6? @ganeshie8

OpenStudy (jacobbirddog):

or 12/6?

ganeshie8 (ganeshie8):

plugin x = 6, y = 10 : \(\large 10 = \frac{k}{6}\) \(\large k = 60\)

OpenStudy (jacobbirddog):

could it be 4.8 or like 7.5?

ganeshie8 (ganeshie8):

so, \(\large y = \frac{60}{x}\) is your inverse function

ganeshie8 (ganeshie8):

read the quesiton, u want to find y when x = 8, right ?

ganeshie8 (ganeshie8):

simply plugin x = 8 above^

ganeshie8 (ganeshie8):

\(\large y = \frac{60}{8} = \frac{15}{2} = 7.5\)

OpenStudy (jacobbirddog):

ok so i was right :D ty @ganeshie8 and thanks for telling me how to di it in detail

ganeshie8 (ganeshie8):

yes :) u wlc !